# Explicit conserved operators for a class of integrable bosonic networks from the classical Yang-Baxter equation

### v1

@@ 1 Introduction
The goal of this work is to combine integrability with bosonic degrees of freedom into the broader picture of networks. This approach is partly motivated by the rising profile of extended Bose-Hubbard models for studies in cold atom systems with dipolar or cavity-mediated interactions [12, 30, 31, 32, 45]. Our investigation below focuses on a class of networks that are associated with bipartite graphs. Examples of this class were introduced in [73] for cases where the associated bipartite graph $K_{p,q}$ is complete. (For an introduction to basic notions in graph theory, see [28].) Particular applications have been made for small-sized systems. Adapting the work of [31], physical realisations of the $K_{2,1}$ system through dipolar atoms have been offered in [61, 68], and several theoretical aspects of the model have since been studied [10, 11, 69, 70, 71]. The $K_{2,2}$ system [60] has found application in the generation of NOON states [7, 23, 24], the $K_{3,1}$ system for state transfer on a quantum turntable [74], and both the $K_{3,2}$ and $K_{4,1}$ systems as means for generating $W$-states [8]. As shown in [6], these models on complete graphs are generally superintegrable, including the star models [19]; they possess more conserved operators than degrees of freedom. This suggests the possibility to relax the completeness property of the graph, whilst maintaining integrability.
@@
where ${\mathcal{B}}$ is a real, symmetric matrix. There are no other constraints placed on ${\mathcal{B}}$. Let $\{a_{j},\,a_{j}^{\dagger}:j=1,\dots,m\}\cup\{b_{j},\,b_{j}^{\dagger}:j=1,\dots,m\}$ denote mutually commuting sets of canonical boson operators satisfying
@@
In Sect. 2 we provide preliminary background to formulate the results on the framework of Lie algebras and dual Poisson structures. Using the classical Yang-Baxter equation allows for the generation of a commutative Poisson subalgebra, representing conservation laws in an abstract dynamical system. This formalism is shown to be realised through the Lie algebra $gl(n)$, and the section concludes with discussion on the problem of quantisation. In Sect. 3 these procedures are implemented for a specific solution of the classical Yang-Baxter equation. It is described how this solution is amenable for the construction of (6) on arbitrary bipartite graphs described above. A route to quantise the commutative Poisson subalgebra is demonstrated. As an application of these results, we establish the integrability of a 4-site Hamiltonian that has been discussed in recent literature [38]. Final comments are gathered in Sect. 5, while an important technical result on the commutativity of transfer matrices is detailed in the Appendix.
@@ 2 Preliminaries
2.1 Lie algebras and their dual Poisson algebras
Here we follow the frameworks set out in [1, 57, 67]. Let $\mathfrak{g}$ denote a Lie algebra with basis $\{X_{j}:j=1,\dots,d\}$ and structure constants $\{c_{jk}^{l}:j,k,l=1,\dots,d\}$ such that the Lie bracket
@@ 2.1 Lie algebras and their dual Poisson algebras
Let $\mathfrak{g}^{*}$ denote the dual algebra of $\mathfrak{g}$, with basis $\{\mathcal{X}_{j}:j=1,\dots,d\}$ such that
$\mathcal{X}_{j}(X_{k})=\delta_{jk}.$
Let $P(\mathfrak{g}^{*})$ denote the associated polynomial algebra generated by $\{\mathcal{X}_{j}:j=1,\dots,d\}$, equipped with a Poisson algebra structure through the bracket
@@
holds. We say that $f,g\in\,P(\mathfrak{g}^{*})$ Poisson-commute if
@@
Let $P_{k}(\mathfrak{g}^{*})$ denote the space of homogeneous polynomials of degree $k$, satisfying
$\{P_{j}(\mathfrak{g}^{*}),\,P_{k}(\mathfrak{g}^{*})\}\subseteq P_{j+k-1}(\mathfrak{g}^{*})$
@@
$\{P_{1}(\mathfrak{g}^{*}),\,P_{k}(\mathfrak{g}^{*})\}\subseteq P_{k}(\mathfrak{g}^{*}).$ (12)
Eq. (9) shows that the Lie algebra $\mathfrak{g}$ is canonically embedded in $P(\mathfrak{g}^{*})$ as $P_{1}(\mathfrak{g}^{*})$, in that the mapping $\mathcal{X}_{j}\mapsto X_{j}$ between basis elements provides a Lie algebra isomorphism. For $\mathcal{Y}_{1},\dots,\mathcal{Y}_{k}\in\,\mathfrak{g}^{*}$ let the corresponding images under this isomorphism be denoted $Y_{1},\dots,Y_{k}\in\,\mathfrak{g}$. Let $S_{k}$ denote the symmetric group on $k$ objects. Define the vector space isomorphism $\iota:P(\mathfrak{g}^{*})\rightarrow U(\mathfrak{g})$ via the following action on products of elements in $\mathfrak{g}^{*}$
@@
and extended linearly to all of $P(\mathfrak{g}^{*})$. Above, $S_{k}$ denote the group of permutations on $k$ objects. Set $U_{k}(\mathfrak{g})=\iota(P_{k}(\mathfrak{g}^{*}))$. It then follows [67]
@@ 2.2 Classical Yang-Baxter equation for generating commutative Poisson subalgebras
Here we follow the framework set out in [25, 50, 51]. Let $V$ denote a complex vector space of dimension $n$, and let $r(u,v)\in\,{\rm End}(V\otimes V)$ with $u,v\in\,{\mathbb{C}}$. Then $r(u,v)$ is termed an $r$-matrix if it is a solution of the classical Yang-Baxter equation
@@
Introduce the set of generating elements $\{\mathcal{T}^{j}_{k}(u):j,k=1,\dots n\}$ for an abstract Poisson algebra ${\mathcal{P}}$ and set
@@
A Poisson bracket is imposed according to
@@
for the Poisson bracket holds as a result of (14).
Define
$(\mathcal{T}^{(1)})^{j}_{k}(u)$ $=\mathcal{T}^{j}_{k}(u),$
$(\mathcal{T}^{(s+1)})^{j}_{k}(u)$ $=\sum_{l=1}^{n}(\mathcal{T}^{(s)})^{j}_{l}(u)\mathcal{T}^{l}_{k}(u)=\sum_{l=1}^{n}\mathcal{T}^{j}_{l}(u)(\mathcal{T}^{(s)})^{l}_{k}(u),$
@@
it follows that $\{\mathcal{I}_{j}^{(r)},\,\mathcal{I}_{k}^{(s)}\}=0$; i.e. these elements form a commutative Poisson subalgebra.
@@ 2.3 Shift elements and realisation of the commutative Poisson subalgebra
$[E^{j}_{k},\,E^{p}_{q}]=\delta^{p}_{k}E^{j}_{q}-\delta^{j}_{q}E^{p}_{k},$
and let $\{\mathcal{E}^{j}_{k}:j,k=1,\dots,n\}$ denote the dual-space basis with Poisson bracket
@@
which we refer to as a shift element following [52]. It is straightforward to check, as a result of (14,19) that the mapping
$\varrho(\mathcal{T}^{j}_{k}(u))=B^{j}_{k}(u)+\sum_{p,q=1}^{n}r^{jp}_{kq}(u,v)\mathcal{E}^{q}_{p}$ (20)
preserves the Poisson bracket relation (15). This realisation of the $\mathcal{T}^{j}_{k}(u)$ in turn affords a realisation to generate a commutative Poisson subalgebra in ${\mathcal{P}}(gl(n)^{*})$.
@@ 2.4 Quantisation
An attempt to quantise the above considerations, by replacing Poisson brackets with commutators, faces a roadblock. Introduce the set of generating elements $\{T^{j}_{k}(u):j,k=1,\dots n\}$ for an abstract classical Yang-Baxter algebra $Y$ and set
@@
As a result of (14) and (21), it is seen that the $r$-matrix affords the defining realisation $\pi_{v}$ of $Y$ through the mapping
${\pi}_{v}(T^{j}_{k}(u))=\sum_{p,q=1}^{n}r^{jp}_{kq}(u,v)E^{q}_{p},$
@@
Symmetries of $r(u,v)$ are identified as those $Z(u)$ for which the skew-symmetric action is identically zero, coinciding with (19). For any one-dimensional realisation $\eta:Y\rightarrow{\mathbb{C}}$ we have
@@
preserves the commutation relations (22), thus providing a one-dimensional realisation. Such one-dimensional realisations encode the symmetries of $r(u,v)$ in a generalised sense, and will play a significant role in later calculations. Finally, taking the tensor product of a one-dimensional realisation with the defining realisation yields a realisation expressible as
@@
Now, following the procedures of Sect. 2.3 further by defining
$(T^{(1)})^{j}_{k}(u)$ $=T^{j}_{k}(u),$
$(T^{(s+1)})^{j}_{k}(u)$ $=\sum_{l=1}^{n}(T^{(s)})^{j}_{l}(u)T^{l}_{k}(u)=\sum_{l=1}^{n}T^{j}_{l}(u)(T^{(s)})^{l}_{k}(u),$
$t^{(s)}(u)$ $=\sum_{j=1}^{n}(\mathcal{T}^{(s)})^{j}_{j}(u)$ (25)
@@
More generally, the problem to quantise elements of a commutative Poisson subalgebra to central elements in the universal enveloping algebra of a Lie algebra has been well-studied with a long history, e.g. [18, 20, 39, 40, 42, 46, 51, 57, 58, 59, 67] A main objective below is to develop techniques suitable for quantising a specific case of commutative, non-homogeneous, polynomial elements in ${\mathcal{P}}(gl(n)^{*})$, where $n$ is even.
@@ 3.1 Solution of the classical Yang-Baxter equation, shift element, and the Hamiltonian
$r(u,v)=\frac{u}{2}\left(\frac{1}{u-v}I\otimes I+\frac{1}{u+v}A\otimes A\right)P.$ (26)
It may be checked directly that (14) is satisfied provided $A^{2}=I$. Next, we turn towards the shift element. Setting $B(u)=uB$ it is found that (19) holds provided $AB=-BA$. These conditions are satisfied by choosing $n=2m$ and
@@
where the canonical boson operators satisfy (5), the identification
@@ 3.2 Commutative Poisson subalgebra
$\varrho_{v}(\mathcal{I}^{(s)}(u))\sim\begin{cases}su^{s-1}\,{\rm tr}(B^{s-1}\mathcal{E}_{\rm e}),&s\,\,{\rm odd},\\ \displaystyle u^{s}\,{\rm tr}(B^{s})+su^{s-2}\left(v\,{\rm tr}(B^{s-1}\mathcal{E}_{\rm o})+\frac{1}{2}\sum_{k=0}^{s-2}{\rm tr}\left(B^{s-2-k}\mathcal{E}_{\rm e}B^{k}\mathcal{E}_{\rm e}\right)\right),&s\,\,{\rm even}.\end{cases}$ (35)
The images under (13) of these elements provide candidates for constructing the conserved operators for the Hamiltonian (6). Although we only take leading-order terms in the expansion above, we will make remarks in the Discussion to argue why these are sufficient for this purpose.
@@ 3.3 Conserved operators from the commutative Poisson subalgebra
By taking the images under (13), recalling that $\mathcal{U}^{-1}=2v$, undertaking rescaling, and streamlining notation leaves us to verify that the following operators are mutually commutative:
@@
Let $\mathcal{C}(2p),\,\mathcal{C}(2p+1),\,\mathcal{D}(2p,i)\in\,{\mathcal{P}}(gl(2m)^{*})$ denote their Poisson algebra counterparts through $E^{a}_{b}\mapsto\mathcal{E}^{a}_{b}$ for all $a,b\in\,{\rm e}\cup{\rm o}$. We have checked, through several pages of long and arduous direct calculations that
@@
it follows that the restricted property $\mathcal{J}\in\,gl(2m)$ must hold. Applying the automorphism $\sigma$ from Subsect. 2.1 shows that $\sigma(\mathcal{J})=\mathcal{J}$, while
@@
Note that there is a canonical embedding $\chi:gl(m)\rightarrow gl(2m)$ given by
@@
Let ${\rm im}(\chi)$ denote the image of $\chi$, which spans the subspace of $gl(2m)$ that is invariant under (34). Hence $\mathcal{K}_{1}\in\,{\rm im}(\chi)$. It is also clear that $C(2p)\in\,{\rm im}(\chi)$, satisfying
@@
We say that the symmetric matrix $\mathcal{B}$ is generic if the eigenvalues of $\mathcal{B}^{2}$ are distinct. In such an instance Eq. (42) indicates that $\{C(2p):p=1,\dots,m\}$ provides a basis for a Cartan subalgebra ${\mathfrak{H}}\subset{\rm im}(\chi)$, Eq. (39) indicates that $\mathcal{K}_{1}$ is in the centraliser of ${\mathfrak{H}}$, and consequently $\mathcal{K}_{1}\in\,{\mathfrak{H}}$ since ${\mathfrak{H}}$ is a maximally commutative subalgebra. In light of (40) and (41) it follows that $\mathcal{K}_{1}=0$. Thus $\mathcal{K}$ as given by (38) is zero. Consequently $[C(j),\,C(k)]=0$ for all $j$ and $k$. The result also holds for non-generic $\mathcal{B}$, since this can be obtained as the limit of a perturbation. Let ${\mathbb{X}}$ denote a matrix that diagonalises $\mathcal{B}$, and ${\mathbb{D}}$ a diagonal matrix chosen such that $\mathcal{B}(\epsilon)=\mathcal{B}+\epsilon{\mathbb{X}}{\mathbb{D}}{\mathbb{X}}^{-1}$ is generic in a neighbourhood of $\mathcal{B}$. Since $[C(j),\,C(k)]=0$ holds for $\mathcal{B}(\epsilon)$, it also holds in the limit $\epsilon\rightarrow 0$ since the $C(j)$ only depend on the matrix elements of the powers of $\mathcal{B}(\epsilon)$.

### v2

@@ 1 Introduction
The goal of this work is to combine integrability with bosonic degrees of freedom into the broader picture of networks. This approach is partly motivated by the rising profile of extended Bose-Hubbard models for studies in cold atom systems with dipolar or cavity-mediated interactions [12, 30, 31, 32, 45]. Our investigation below focuses on a class of networks that are associated with bipartite graphs. Examples of this class were introduced in [77] for cases where the associated bipartite graph $K_{p,q}$ is complete. (For an introduction to basic notions in graph theory, see [28].) Particular applications have been made for small-sized systems. Adapting the work of [31], physical proposals for the $K_{2,1}$ system through dipolar atoms have been offered in [65, 72], and several theoretical aspects of the model have since been studied [10, 11, 73, 74, 75]. The $K_{2,2}$ system [64] has found application in the generation of NOON states [7, 23, 24], the $K_{3,1}$ system for state transfer on a quantum turntable [78], and both the $K_{3,2}$ and $K_{4,1}$ systems as means for generating $W$-states [8]. As shown in [6], these models on complete bipartite graphs are generally superintegrable, including the star models [19]; they possess more conserved operators than degrees of freedom. This suggests the possibility to relax the completeness property of the bipartite graph, whilst maintaining integrability.
@@
where $\mathcal{B}$} is a real, symmetric matrix. There are no other constraints placed on $\mathcal{B}$}. Let $\{a_{j},\,a_{j}^{\dagger}:j=1,\dots,m\}\cup\{b_{j},\,b_{j}^{\dagger}:j=1,\dots,m\}$ denote mutually commuting sets of boson operators satisfying
@@
In Sect. 2 we provide preliminary background to formulate the results on the framework of Lie algebras and Poisson structures. Using the classical Yang-Baxter equation allows for the generation of a commutative Poisson subalgebra, representing conservation laws in an abstract dynamical system. This formalism is shown to be realised through the Lie algebra $gl(n)$, and the section concludes with discussion on the problem of quantisation. In Sect. 3 these procedures are implemented for a specific solution of the classical Yang-Baxter equation. It is described how this solution is amenable for the construction of (6) on arbitrary bipartite graphs described above. A route to quantise the commutative Poisson subalgebra is demonstrated. As an application of these results, we establish the integrability of a 4-site Hamiltonian that has been discussed in recent literature [38]. Final comments are gathered in Sect. 5, while an important technical result on the commutativity of transfer matrices is detailed in the Appendix.
@@ 2 Preliminaries
2.1 Lie algebras and related Poisson algebras
Here we follow the frameworks set out in [1, 61, 71]. Let $\mathfrak{g}$ denote a finite-dimensional, complex Lie algebra with basis $\{X_{j}:j=1,\dots,d\}$ and structure constants $\{c_{jk}^{l}:j,k,l=1,\dots,d\}$ such that the Lie bracket
@@ 2.1 Lie algebras and related Poisson algebras
Definition 1
Let ${\mathbb{C}}[\mathcal{X}_{1},\dots,\mathcal{X}_{d}]$ denote the ring of polynomial functions in $d$ indeterminates. Let ${\mathcal{A}}(\mathfrak{g})$ denote the algebra obtained by defining a Poisson structure on ${\mathbb{C}}[\mathcal{X}_{1},\dots,\mathcal{X}_{d}]$ through the bracket
@@
holds. We say that $f,g\in\,{\mathcal{A}}(\mathfrak{g})$ Poisson-commute if
@@
Let ${\mathcal{A}}_{k}(\mathfrak{g})$ denote the space of homogeneous polynomials of degree $k$, satisfying
$\{{\mathcal{A}}_{j}(\mathfrak{g}),\,{\mathcal{A}}_{k}(\mathfrak{g})\}\subseteq{\mathcal{A}}_{j+k-1}(\mathfrak{g})$
@@
$\{{\mathcal{A}}_{1}(\mathfrak{g}),\,{\mathcal{A}}_{k}(\mathfrak{g})\}\subseteq{\mathcal{A}}_{k}(\mathfrak{g}).$ (12)
Eq. (9) shows that the Lie algebra $\mathfrak{g}$ is embedded in ${\mathcal{A}}(\mathfrak{g})$ as ${\mathcal{A}}_{1}(\mathfrak{g})$, in that the mapping $\mathcal{X}_{j}\mapsto X_{j}$ between basis elements provides a Lie algebra isomorphism. For $\mathcal{Y}_{1},\dots,\mathcal{Y}_{k}\in\,{\mathcal{A}}_{1}(\mathfrak{g})$ let the corresponding images under this isomorphism be denoted $Y_{1},\dots,Y_{k}\in\,\mathfrak{g}$. Let $S_{k}$ denote the symmetric group on $k$ objects. Define the vector space isomorphism $\iota:{\mathcal{A}}(\mathfrak{g})\rightarrow U(\mathfrak{g})$ via the following action on products of elements in ${\mathcal{A}}_{1}(\mathfrak{g})$
@@
and extended linearly to all of ${\mathcal{A}}(\mathfrak{g})$. Set $U_{k}(\mathfrak{g})=\iota({\mathcal{A}}_{k}(\mathfrak{g}))$. It then follows [71]
@@ 2.2 Classical Yang-Baxter equation for generating commutative Poisson subalgebras
Here we follow the framework set out in [25, 50, 52]. Let $V$ denote a complex vector space of dimension $n$, and let $r(u,v)\in\,{\rm End}(V\otimes V)$ with $u,v\in\,{\mathbb{C}}$. Then $r(u,v)$ is termed an $r$-matrix if it is a solution of the classical Yang-Baxter equation [3, 4, 27, 29, 47, 49, 52, 57]
@@
Definition 2
Given a solution $r(u,v)$ of the classical Yang-Baxter equation (15), introduce the set of generating elements $\{\mathcal{T}^{j}_{k}(u):j,k=1,\dots n\}$, for an abstract Poisson algebra $\mathcal{P}$}. Here, $u$ is a formal variable such that
$\mathcal{T}^{j}_{k}(u)=\sum_{p=-\infty}^{\infty}[\mathcal{T}^{j}_{k}]_{p}\,u^{p}.$
Set
@@
The Poisson bracket is imposed according to
@@
for the Poisson bracket holds as a result of $r(u,v)$ satisfying the classical Yang-Baxter equation (15).
Define the elements $(\mathcal{T}^{(s)})^{j}_{k}(u)$ through powers of the generating tensor
$[\mathcal{T}(u)]^{s}=\sum_{j,k=1}^{n}e^{k}_{j}\otimes(\mathcal{T}^{(s)})^{j}_{k}(u)$
and set
@@
it follows that
$\{\mathcal{I}_{j}^{(r)},\,\mathcal{I}_{k}^{(s)}\}=0,$ (17)
i.e. these elements form a commutative Poisson subalgebra.
@@ 2.3 Shift elements and realisation of the commutative Poisson subalgebra
$[E^{j}_{k},\,E^{p}_{q}]=\delta^{p}_{k}E^{j}_{q}-\delta^{j}_{q}E^{p}_{k}$
@@
Let $\{\mathcal{E}^{j}_{k}:j,k=1,\dots,n\}$ denote indeterminates with Poisson bracket
@@
which we refer to as a shift element following [53]. It is straightforward to check, as a result of (15,20) that the mapping $\varrho_{v}:{\mathcal{P}}\rightarrow{\mathcal{A}}(gl(n))$ defined by
$\varrho_{v}(\mathcal{T}^{j}_{k}(u))=B^{j}_{k}(u)+\sum_{p,q=1}^{n}r^{jp}_{kq}(u,v)\mathcal{E}^{q}_{p}$ (21)
preserves the Poisson bracket relation (16). We refer to $\varrho_{v}$ as the realisation of $\mathcal{P}$} afforded by ${\mathcal{A}}(gl(n))$. This realisation provides the means to generate a commutative Poisson subalgebra in ${\mathcal{A}}(gl(n))$ through the images of the elements ${\mathcal{I}}^{(s)}_{k}$ under $\varrho_{v}$.
@@ 2.4 Quantisation
An attempt to quantise the above considerations, by replacing Poisson brackets with commutators, faces a roadblock, as will be explained below.
Definition 3
Introduce the set of generating elements $\{T^{j}_{k}(u):j,k=1,\dots n\}$, where $u$ denotes a formal variable, for an abstract classical Yang-Baxter algebra $Y$. Set
@@
As a result of (15) and (22), it is seen that the $r$-matrix affords the defining realisation $\pi_{v}:Y\rightarrow U(gl(n))$ through the homomorphism
${\pi}_{v}(T^{j}_{k}(u))=\sum_{p,q=1}^{n}r^{jp}_{kq}(u,v)E^{q}_{p}.$
@@
Symmetries of $r(u,v)$ are identified as those $Z(u)$ for which the skew-symmetric action is identically zero, coinciding with (20). For any one-dimensional representation $\eta:Y\rightarrow{\mathbb{C}}$ we have
@@
preserves the commutation relations (23), thus providing a one-dimensional representation. Such one-dimensional representations encode the symmetries of $r(u,v)$ in a generalised sense, and will play a significant role in later calculations. Finally, taking the tensor product of a one-dimensional representation with the defining realisation yields a realisation expressible as
@@
Now, following the procedures of Sect. 2.3 further by defining $(T^{(s)})^{j}_{k}(u)$ through powers of the generating tensor
$[T(u)]^{s}=\sum_{j,k=1}^{n}e^{k}_{j}\otimes(T^{(s)})^{j}_{k}(u)$
and setting
$t^{(s)}(u)$ $=\sum_{j=1}^{n}(T^{(s)})^{j}_{j}(u)$ (26)
@@
More generally, the problem to quantise elements of a commutative Poisson subalgebra to an abelian subalgebra in the universal enveloping algebra of a Lie algebra has been well-studied with a long history, e.g. [18, 20, 39, 40, 42, 46, 52, 61, 62, 63, 71] A main objective below is to develop techniques suitable for quantising a specific case of commutative, non-homogeneous, polynomial elements in ${\mathcal{A}}(gl(n))$, where $n$ is even.
@@ 3.1 Solution of the classical Yang-Baxter equation, shift element, and the Hamiltonian
$r(u,v)=\frac{u}{2}\left(\frac{1}{u-v}I\otimes I+\frac{1}{u+v}A\otimes A\right)P,$ (27)
where $A\in{\rm End}({\mathbb{C}}^{n})$. It may be checked directly that (15) is satisfied provided $A^{2}=I$. This solution belongs to a class discussed in [51]. Next, we turn towards obtaining a non-constant shift element, c.f. [53]. Setting $B(u)=uB$ it is found that (20) holds provided AB=BAAB=-BA. These conditions are satisfied by choosing $n=2m$ and
@@
where the boson operators satisfy (5), the identification
@@ 3.2 Commutative Poisson subalgebra
$\varrho_{v}(\mathcal{I}^{(s)}(u))\sim\begin{cases}su^{s-1}\,{\rm tr}(B^{s-1}\mathcal{E}_{\rm e}),&s\,\,{\rm odd},\\ \displaystyle u^{s}\,{\rm tr}(B^{s})+su^{s-2}\left(v\,{\rm tr}(B^{s-1}\mathcal{E}_{\rm o})+\frac{1}{2}\sum_{k=0}^{s-2}{\rm tr}\left(B^{s-2-k}\mathcal{E}_{\rm e}B^{k}\mathcal{E}_{\rm e}\right)\right),&s\,\,{\rm even}.\end{cases}$
and in particular
$\varrho_{v}(\mathcal{I}_{s-1}^{(s)}(u))$ $=s\,{\rm tr}(B^{s-1}\mathcal{E}_{\rm e}),$ $s\,\,{\rm odd},$ (36)
$\varrho_{v}(\mathcal{I}_{s-2}^{(s)}(u))$ $=s\left(v\,{\rm tr}(B^{s-1}\mathcal{E}_{\rm o})+\frac{1}{2}\sum_{k=0}^{s-2}{\rm tr}\left(B^{s-2-k}\mathcal{E}_{\rm e}B^{k}\mathcal{E}_{\rm e}\right)\right),$ $s\,\,{\rm even}.$
The images of (36) under (13) provide candidates for constructing the conserved operators for the Hamiltonian (6). Although we only take leading-order terms in the expansion above, we will make remarks in the Discussion to argue why these are sufficient for this purpose.
@@ 3.3 Conserved operators from the commutative Poisson subalgebra
By taking the images of (36) under (13), recalling that $\mathcal{U}^{-1}=2v$, undertaking rescaling, and streamlining notation leaves us to verify that the following operators are mutually commutative:
@@
Let $\mathcal{C}(2p),\,\mathcal{C}(2p+1),\,\mathcal{D}(2p,i)\in\,{\mathcal{A}}(gl(2m))$ denote their Poisson algebra counterparts through $E^{a}_{b}\mapsto\mathcal{E}^{a}_{b}$ for all $a,b\in\,{\rm e}\cup{\rm o}$. We have checked, through several pages of long and arduous direct calculations that
@@
it follows that the restricted property $\mathcal{J}\in\,gl(2m)$ must hold. Applying the automorphism $\sigma$ given by (14) shows that $\sigma(\mathcal{J})=\mathcal{J}$, while
@@
Note that there is an embedding $\chi:gl(m)\rightarrow gl(2m)$ given by
@@
Let ${\rm im}(\chi)\cong gl(m)$ denote the image of $\chi$, which spans the subspace of $gl(2m)$ that is invariant under (35). Hence $\mathcal{K}_{1}\in\,{\rm im}(\chi)$. It is also clear that $C(2p)\in\,{\rm im}(\chi)$, satisfying
@@
We say that the symmetric matrix $\mathcal{B}$ is generic if the eigenvalues of $\mathcal{B}^{2}$ are distinct. In such an instance Eq. (43) indicates that $\{C(2p):p=1,\dots,m\}$ provides a basis for a Cartan subalgebra ${\mathfrak{H}}\subset{\rm im}(\chi)$, Eq. (40) indicates that $\mathcal{K}_{1}$ is in the centraliser of $\mathfrak{H}$}, and consequently $\mathcal{K}_{1}\in\,{\mathfrak{H}}$ since $\mathfrak{H}$} is a maximally commutative subalgebra. In light of (41) and (42) it follows that $\mathcal{K}_{1}=0$. Thus $\mathcal{K}$ as given by (39) is zero. Consequently [C(j),C(k)]=0[C(j),\,C(k)]=0 for all $j$ and $k$. The result also holds for non-generic $\mathcal{B}$, since this can be obtained as the limit of a perturbation. Let $\mathbb{X}$} denote a matrix that diagonalises $\mathcal{B}$, and $\mathbb{D}$} a diagonal matrix chosen such that $\mathcal{B}(\epsilon)=\mathcal{B}+\epsilon{\mathbb{X}}{\mathbb{D}}{\mathbb{X}}^{-1}$ is generic in a neighbourhood of $\mathcal{B}$. Since [C(j),C(k)]=0[C(j),\,C(k)]=0 holds for $\mathcal{B}(\epsilon)$, it also holds in the limit $\epsilon\rightarrow 0$ since the $C(j)$ only depend on the matrix elements of the powers of $\mathcal{B}(\epsilon)$.